OpenStudy (anonymous):

If you were going to eliminate the x's in the system below, you would multiply: 6x + 2y = 2 4x - 3y = 11 A. the top equation by 6, the bottom by 2 B. the top equation by 6, the bottom by 4 C. the top equation by -2, the bottom by 3 D. the top equation by 3, the bottom by 2

OpenStudy (lastdaywork):

What's your guess ??

OpenStudy (anonymous):

I think A

OpenStudy (anonymous):

@LastDayWork

OpenStudy (lastdaywork):

According to Option A 6*(6x + 2y) = 6*2 36x + 12y = 12 2*(4x - 3y) = 2*11 8x - 6y = 22 Do you still think A is the correct ??

OpenStudy (anonymous):

nope B

OpenStudy (lastdaywork):

Do a similar analysis for all the options (like I did for A) and then tell me the answer..